# Where Quantum Complexity Helps Classical Complexity: A Survey

### v2

@@ 4.16. Knapsack Problem
García and Maureira (garcia2021knn, ) introduced a hybrid algorithm, using the k-nearest neighbor technique, improving over a quantum cuckoo search algorithm applied in resource allocation. The authors performed numerical experiments, obtaining insights from the involvement of the k-nearest neighbor technique in their final result. The multidimensional knapsack problem was explored in (garcia2021knn, ) as well, for the validation of their procedure. Further, the authors made a comparison to SOTA algorithms. They illustrated that their hybrid algorithm offers better results in most cases.
Lai et al. (lai2020diversity, ) have explored a quantum particle swarm optimization algorithm, integrating a diversity-preserving strategy for population management. Also, they included a local optimization method based on variable neighborhood descent, improving solutions. The method in (lai2020diversity, ) was evaluated on the classic NP-hard 0–1 multidimensional knapsack problem (MKP). The experimental results, obtained by the authors, on $270$ instances showed that their algorithm is efficient compared to several SOTA MKP algorithms. This supremacy holds in two terms: the solution quality and the computational efficiency on the instances with a small or medium-sized number $(m\leq 10)$ of constraints.
@@ 4.17. Subset-Sum Problem
Helm and May (helm2020power, ) proposed quantum subset sum algorithms, not using QRAM, and breaking the square root Grover bound using linear many qubits. In particular, they obtain a running time of $2^{0.48n}$. This result is built upon the representation technique and the quantum collision finding algorithm, suggested by Chailloux, Naya-Plasencia, and Schrottenloher.
The authors in (helm2020power, ) nonetheless improve the above result to $2^{0.43n}$. The critical point about the trade-off they introduce requires using classical memory only. Namely, not QRAM. There is a time-memory-qubit trade-off. Their result is obtained by using the Schroeppel-Shamir list construction technique. The second algorithm needs to be compared to purely classical time-memory subset sum trade-offs. For instance, those of Howgrave-Graham and Joux. In particular, the algorithms of Helm and May improve on the above completely classical algorithms for all memory complexities less than $2^{0.2n}$.
@@ 4.18. Coding and Cryptography
Hhan et al. (hhan2023quantum, ) studied the quantum complexity of discrete logarithms and related group-theoretic problems. The authors considered their settings under generic algorithms that do not exploit the properties of the group encoding. They set up a generic model of quantum computation named the quantum generic group model. This model is the quantum variant of its classical counterpart. The authors pointed out that one may describe Shor’s algorithm for the discrete logarithm problem and related algorithms in their suggested model. Under their model, Hhan et al. showed quantum lower bounds, also (almost) matching algorithms of the discrete logarithm and related problems In particular, consider a cyclic group $\mathcal{G}$ of prime order. The authors proved that (*)(*)( * ) the depth of group operation queries made by every generic quantum discrete logarithm algorithm is lower bounded by $\Omega(\log\mathcal{G})$. This result implies that Shor’s algorithm is asymptotically optimal among the generic quantum algorithms, even under parallel settings. The authors further observed that (**)(**)( * * ) some known variations of Shor’s algorithm may benefit from classical computations to reduce the number and depth of quantum group operations. They suggested a model for generic hybrid quantum-classical algorithms. It shows that these algorithms are almost optimal in this model. All generic hybrid quantum-classical algorithms for the discrete logarithm problem with a total number of (classical or quantum) group operations $Q$ must make $\Omega(\log|G|/\log Q)$ quantum group operations of depth $\Omega(\log\log|G|-\log\log Q)$. Specifically, if $Q=\text{polylog}|G|$, classical group operations reduce the number of quantum queries by a factor of $\mathcal{O}(\log\log|G|)$. And, the quantum depth remains as $\Omega(\log\log|G|)$.
@@
Aaronson and Hung (aaronson2023certified, ) suggested an application for near-term quantum devices. That is, constructing cryptographically certified random bits so that one may use them in proof-of-stake cryptocurrencies, as well as in other areas. The protocol adapts the currently available quantum supremacy experiments. This would be based on random circuit sampling. Let $n$ be the number of qubits. The authors particularly showed that under plausible hardness assumptions, every time the outputs of the latter experiments pass the now-standard Linear Cross-Entropy Benchmark, they necessarily contain $\Omega(n)$ min-entropy. Aaronson and Hung (aaronson2023certified, ) obtained a net gain in randomness using a small seed to generate pseudorandom challenge circuits. To have a response in return for the challenge, the quantum computer has to generate output strings. The strings have to be verified first and then one may provide them as inputs into a randomness extractor. This will produce certified nearly-uniform bits. Hence bootstrapping from pseudorandomness to genuine randomness. The authors show that their protocol reasonably works. This holds from two points of view. First, under a hardness assumption named Long List Quantum Supremacy Verification. In particular, they gave ground for it in the random oracle model. Second, they considered an eavesdropper who could share arbitrary entanglement with the device. The authors’ proof was unconditionally in the random oracle model against the aforementioned eavesdropper. One should note that the output of the protocol is unpredictable. Even to an unbounded adversary in terms of computational power, who may see the random oracle.
A drawback of the above protocol is the exponential cost of verification. Specifically, there cannot be more than $n\sim 60$ qubits. Namely, a setting in which attacks are costly, yet not impractical.
Leverrier and Zémor (leverrier2023efficient, ) introduced and analyzed a decoder for quantum tanner codes. These codes may correct adversarial errors of linear weight. Notice that the previous decoders for quantum low-density parity-check codes supported weights of $\mathcal{O}(\sqrt{n\log n})$ (ShaiEvra2020, ).
Further, the authors in (leverrier2023efficient, ) contributed to the link between quantum tanner codes and the Lifted Product codes of Panteleev and Kalachev (panteleev2022asymptotically, ). They proved that one may adapt their decoder to the latter. The decoding algorithm in (leverrier2023efficient, ) alternates between sequential and parallel procedures. It also offers convergence in linear time.
@@
5. Open Problems
@@ 5.1. QAOA
Problem 1 ().
Adaptive-bias QAOA (ab-QAOA) is used to improve performance in the hard region of the 3-SAT and the Max-3-SAT problems (yu2023solution, ). Initialization of bias-field parameters $\vec{h}$, which can be generated randomly, is a key issue. Consider the way to obtain the same infidelity using a very large computational resource. This is a matter of understanding how accurate the initial guess on $\vec{h}$ scales with $n$. Is it then possible to take a non-random approach, using a classical heuristic, to get the initial value of $\vec{h}$?
Problem 2 ().
It is known that the gap between the QAOA and the abQAOA (adaptive-bias QAOA) is greatest when the plateau barrens  [^3] The barren plateau phenomenon explains that the gradient of parameters of the parameterized quantum circuits will vanish exponentially in terms of the system size (Zhao_2021, ) are most visible. Equally, numerical evidence suggests that abQAOA has better scaling behavior than QAOA in terms of system size accuracy (See (yu2023solution, ) for more information). Is it possible to study a barren plateau phenomenon with abQAOA?
Problem 3 ().
@@ 5.3. Quantum Search
Problem 15 ().
There can be improvements to the running time of the SAT problem by bounding an area and keeping the focus on it in the search space. Making the selected area smaller results in a better running time. Nonetheless, no solution might be found because of a not-right-selected area. Are there approaches to finding an optimal reduction and selection? Even more interestingly, for the larger-sized problems, there may be several areas that may lead to an answer. Can all such areas be found to guarantee the algorithm returns a solution if exists? Moreover, is it possible to adapt the techniques discussed in (varmantchaonala2023quantum, ) to the XORSAT problem?
Problem 16 ().
@@
Problem 17 ().
The work in (CHENG2007123, ) proposed a quantum cooperative search that makes Grover’s search algorithm work on 3-SAT problems with a small number of qubits. Is it possible to conduct the quantum cooperative search (QCS) algorithm to handle other combinatorial optimization problems than 3-SAT?
Problem 18 ().
@@ 5.6. Qubit Mapping
Problem 33 ().
The feasibility of 2048 sparsely connected qubits architectures has already been explored in the literature, for solving SAT and MaxSAT problems as QA systems scale (bian2020solving, ). The existing techniques allow encoding SAT and MaxSAT, with certain limitations, in an Ising problem compatible with sparse QA architecture. Is it possible to use more connected topologies? One may have larger QA hardware graphs with higher per-qubit connectivity and less separation between clusters (tiles) of qubits. New encoding strategies might be required as well because these changes will result in the ability to solve more difficult Ising problems. The encoding process will significantly improve by taking advantage of additional connectivity using new methods for problem decomposition, placing small Boolean functions, penalty modeling, etc.
Problem 34 ().
@@ 5.9. Quantum Merging Trees
5.10. Quantum Congested Clique
Problem 47 ().
@@ 5.11. Verification and Certification
Problem 51 ().
[…]
Problem 52 ().
@@
5.12. Machine Learning
Problem 57 ().
A hybrid algorithm already exists to solve medium and large instances of the multidimensional knapsack problem (garcia2021knn, ). Is it possible to explore the use of dynamic parameters instead of fixed parameters defined in (garcia2021knn, ) for the optimization process? The dynamic parameters might be capable of adjusting as the algorithm is executed, resulting in more powerful algorithms. Do reinforcement learning techniques help to adjust the parameters?
Problem 58 ().
@@ 5.13. Complexity
Problem 63 ().
@@
[…]
are all separations of hybrid quantum depth, and are all according to search problems. Can the same separations be proved w.r.t. decision problems?
Problem 64 ().
Problem 65 ().
[…]
Problem 66 ().
Problem 67 ().
Problem 68 ().
Problem 69 ().
Problem 70 ().
Problem 71 ().
Problem 72 ().
[…]
Problem 73 ().
[…]
and
[…]
@@
Problem 76 ().
[…]
Problem 77 ().
@@ 5.14. Classification
[…]
@@
Problem 83 ().
Is there a way to find the smallest graph (and the smallest number of colors) exhibiting a separation between classical and quantum chromatic numbers?
Problem 84 ().

### v4

@@ 5.16. Knapsack Problem
García and Maureira (garcia2021knn, ) introduced a hybrid algorithm, using the k-nearest neighbor technique, improving over a quantum cuckoo search algorithm applied in resource allocation. The authors performed numerical experiments, obtaining insights from the involvement of the k-nearest neighbour technique in their final result. The multidimensional knapsack problem was explored in (garcia2021knn, ) as well, for the validation of their procedure. Further, the authors made a comparison to SOTA algorithms. They illustrated that their hybrid algorithm offers better results in most cases.
Lai et al. (lai2020diversity, ) have explored a quantum particle swarm optimization algorithm, integrating a diversity-preserving strategy for population management. Also, they included a local optimization method based on variable neighbourhood descent, improving solutions. The method in (lai2020diversity, ) was evaluated on the classic NP-hard 0–1 multidimensional knapsack problem (MKP). The experimental results, obtained by the authors, on $270$ instances showed that their algorithm is efficient compared to several SOTA MKP algorithms. This supremacy holds in two terms: the solution quality and the computational efficiency on the instances with a small or medium-sized number $(m\leq 10)$ of constraints.
@@ 5.17. Subset-Sum Problem
Helm and May (helm2020power, ) proposed quantum subset sum algorithms, not using QRAM, and breaking the square root Grover bound using linear many qubits. In particular, they obtain a running time of $2^{0.48n}$. This result is built upon the representation technique and the quantum collision finding algorithm, suggested by Chailloux, Naya-Plasencia, and Schrottenloher. Helm and May, nonetheless, improved the above result to $2^{0.43n}$. The critical point about the trade-off they introduce requires using classical memory only. Namely, not QRAM. There is a time-memory-qubit trade-off. Their result is obtained by using the Schroeppel-Shamir list construction technique. The second algorithm needs to be compared to purely classical time-memory subset sum trade-offs. For instance, those of Howgrave-Graham and Joux. In particular, the algorithms of Helm and May improve on the above completely classical algorithms for all memory complexities less than $2^{0.2n}$.
@@ 5.18. Coding and Cryptography
Definition 5.25 (Discrete Logarithm).
For every prime number $p$, the multiplicative group $(\mod p)$ forms a cyclic group. Namely, there exist generators $g$ such that $1,\ g,\ g^{2},\ \cdots,\ g^{p-2}$ cover all the non-zero residues (hardy1979introduction, ; knuth1981seminumerical, ) $(\mod p)$. Given a prime number $p$ and such a generator $g$, the discrete logarithm of a number $x$ with respect to $p$ and $g$ is the integer $r$ such that $10\leq r\leq p-$ 1 and $g^{r}\equiv x(\mod p)$ (shor1999polynomial, ).
Classically, there are algorithms that can solve the Discrete Logarithm problem in polynomial time using quantum computation techniques (shor1999polynomial, ; hales2000improved, ; mosca2004exact, ).
Ekerå and Håstad (ekeraa2017quantum, ) extended on Ekerå’s work (ekeraa2016modifying, ) to allow for different trade-offs between the number of times the algorithm needs to be executed, its complexity, and the requirements it imposes on the quantum computer. Their algorithm computes the short discrete logarithm with only a polynomial number of executions. By using lattice-based techniques in a classical post-processing stage, the algorithm provides options for balancing execution frequency and computational complexity. And, it does not require explicit knowledge of the group’s order. Further, the algorithm can solve factoring RSA integers with lower quantum resource requirements.
The work of Ekerå and Håstad (ekeraa2017quantum, ) has been revisited by Ekerå (ekeraa2020post, ), revealing a higher success probability than previously reported. Building on this improved understanding, an enhanced post-processing algorithm has been proposed. This new algorithm allows for better tradeoffs and requires fewer runs. To support these claims, Ekerå developed a classical simulator to sample the probability distribution induced by the quantum algorithm for given logarithms. This simulator represents a significant contribution and is instrumental in demonstrating that the work of Ekerå and Håstad offers an advantage over Shor, not only in individual runs but also overall, particularly when targeting cryptographically relevant instances of RSA and Diffie–Hellman with short exponents.
Roetteler et al. (roetteler2017quantum, ) provided detailed resource estimates for implementing Shor’s algorithm, including the number of qubits and Toffoli gates required for elliptic curve point operations (ECC). They developed reversible algorithms for modular quantum arithmetic, such as addition, subtraction, and multiplication, specifically tailored for ECC. The authors utilized the Weierstrass model to refer to the elliptic curves and elaborated on the group law for point addition. Additionally, they delved into the computational complexity of the Elliptic Curve Discrete Logarithm problem and explored the impact of quantum algorithms on the security of ECC.
The research by Roetteler et al. (roetteler2017quantum, ) was carried forward by Häner et al. (haner2020improved, ). They introduced enhanced quantum circuits designed for elliptic curve scalar multiplication in elliptic curve groups. Their approach involved optimizing reversible integer and modular arithmetic components using windowing techniques and refining the placement of computing steps. Additionally, they enhanced previous quantum circuits for modular inversion by reconfiguring the binary Euclidean algorithm. They developed an affine Weierstrass point addition circuit with reduced depth and fewer $T$ gates compared to earlier circuits. While previous studies focused mainly on minimizing the total number of qubits, this research presents different trade-offs across various cost metrics, including the number of qubits, circuit depth, and T-gate count.
@@
Hhan et al. (hhan2023quantum, ) studied the quantum complexity of discrete logarithms and related group-theoretic problems. The authors considered their settings under generic algorithms that do not exploit the properties of the group encoding. They set up a generic model of quantum computation named the quantum generic group model. This model is the quantum variant of its classical counterpart. The authors pointed out that one may describe Shor’s algorithm for the discrete logarithm problem and related algorithms in their suggested model. Under their model, Hhan et al. showed quantum lower bounds, also (almost) matching algorithms of the discrete logarithm and related problems In particular, consider a cyclic group $\mathcal{G}$ of prime order. The authors proved that the depth of group operation queries made by every generic quantum discrete logarithm algorithm is lower bounded by $\Omega(\log\mathcal{G})$. This result implies that Shor’s algorithm is asymptotically optimal among the generic quantum algorithms, even under parallel settings. The authors further observed that some known variations of Shor’s algorithm may benefit from classical computations to reduce the number and depth of quantum group operations. They suggested a model for generic hybrid quantum-classical algorithms. It shows that these algorithms are almost optimal in this model. All generic hybrid quantum-classical algorithms for the discrete logarithm problem with a total number of (classical or quantum) group operations $Q$ must make $\Omega(\log|G|/\log Q)$ quantum group operations of depth $\Omega(\log\log|G|-\log\log Q)$. Specifically, if $Q=\text{polylog}|G|$, classical group operations reduce the number of quantum queries by a factor of $\mathcal{O}(\log\log|G|)$. And, the quantum depth remains as $\Omega(\log\log|G|)$.
@@
Aaronson and Hung (aaronson2023certified, ) suggested an application for near-term quantum devices. That is, constructing cryptographically certified random bits so that one may use them in proof-of-stake cryptocurrencies, as well as in other areas. The protocol adapts the currently available quantum supremacy experiments. This would be based on random circuit sampling. Let $n$ be the number of qubits. The authors particularly showed that under plausible hardness assumptions, every time the outputs of the latter experiments pass the now-standard Linear Cross-Entropy Benchmark, they necessarily contain $\Omega(n)$ min-entropy. Aaronson and Hung (aaronson2023certified, ) obtained a net gain in randomness using a small seed to generate pseudorandom challenge circuits. To have a response in return for the challenge, the quantum computer has to generate output strings. The strings have to be verified first and then one may provide them as inputs into a randomness extractor. This will produce certified nearly-uniform bits. Hence bootstrapping from pseudorandomness to genuine randomness. The authors showed that their protocol reasonably works. This holds from two points of view. First, under a hardness assumption named Long List Quantum Supremacy Verification. In particular, they gave ground for it in the random oracle model. Second, they considered an eavesdropper who could share arbitrary entanglement with the device. The authors’ proof was unconditionally in the random oracle model against the aforementioned eavesdropper. One should note that the output of the protocol is unpredictable. Even to an unbounded adversary in terms of computational power, who may see the random oracle. A drawback of the above protocol is the exponential cost of verification. Specifically, there cannot be more than $n\sim 60$ qubits. Namely, a setting in which attacks are costly, yet not impractical.
Leverrier and Zémor (leverrier2023efficient, ) introduced and analyzed a decoder for quantum tanner codes. These codes may correct adversarial errors of linear weight. Notice that the previous decoders for quantum low-density parity-check codes supported weights of $\mathcal{O}(\sqrt{n\log n})$ (ShaiEvra2020, ). Further, Leverrier and Zémor contributed to the link between quantum tanner codes and the Lifted Product codes of Panteleev and Kalachev (panteleev2022asymptotically, ). They proved that one may adapt their decoder to the latter. The decoding algorithm in (leverrier2023efficient, ) alternates between sequential and parallel procedures. It also offers convergence in linear time.
@@
6. Open Problems in Applying Quantum Techniqes
@@ 6.1. QAOA
Problem 1.
Adaptive-bias QAOA (ab-QAOA) is used to improve performance in the hard region of the 3-SAT and the Max-3-SAT problems (yu2023solution, ). Initialization of bias-field parameters $\vec{h}$, which can be generated randomly, is a key issue. Consider the way to obtain the same infidelity using an extensive computational resource. This is a matter of understanding how accurate the initial guess on $\vec{h}$ scales with $n$. Is it then possible to take a non-random approach, using a classical heuristic, to get the initial value of $\vec{h}$?
Problem 2.
It is known that the gap between the QAOA and the abQAOA (adaptive-bias QAOA) is greatest when the plateau barrens  [^3] The barren plateau phenomenon explains that the gradient of parameters of the parameterized quantum circuits will vanish exponentially in terms of the system size (Zhao_2021, ) are most visible. Equally, numerical evidence suggests that abQAOA has a better scaling behavior than QAOA in terms of system size accuracy (See (yu2023solution, ) for more information). Is it possible to study a barren plateau phenomenon with abQAOA?
Problem 3.
Recursive quantum approximate optimization algorithm (RQAOA) could be a useful application for NISQ devices in combination with effective classical computation that it may apply to large problems of global relevance (See (bravyi2022hybrid, )). The potential of RQAOA at higher levels $P$ has already been shown, which might not be simulant classical. Is it then possible to develop performance guarantees for $\text{RQAOA}_{p}$, comparable to the robust results achieved by current classical algorithms? For instance, it was previously shown that for the ring of disagrees, RQAOA achieves the optimal approximation ratio. Are there more promising classes of graphs that could yield achievability results?
Problem 4.
It seems beneficial to apply NISQ devices combined with classical computation for recursive quantum approximate optimization algorithm (RQAOA) (bravyi2022hybrid, ). Is it then possible to explore establishing performance guarantees for $\text{RQAOA}_{p}$? Can the next step be finding more classes of graphs s.t. achievability results can be established for them?
Problem 5.
Problem 6.
@@ 6.3. Quantum Search
Problem 15.
There can be improvements to the running time of the SAT problem by bounding an area and keeping the focus on it in the search space. Making the selected area smaller results in a better running time. Nonetheless, no solution might be found because of a not-right-selected area. Are there approaches to finding an optimal reduction and selection? Even more interestingly, for the larger problems, several areas may lead to an answer. Can all such areas be found to guarantee the algorithm returns a solution if exists? Moreover, is it possible to adapt the techniques discussed in (varmantchaonala2023quantum, ) to the XORSAT problem?
Problem 16.
@@
Problem 17.
The work in (CHENG2007123, ) proposed a quantum cooperative search that makes Grover’s search algorithm work on 3-SAT problems with a small number of qubits. Can the quantum cooperative search (QCS) algorithm be conducted to handle combinatorial optimisation problems other than 3-SAT?
Problem 18.
@@ 6.6. Qubit Mapping
Problem 32.
The feasibility of 2048 sparsely connected qubits architectures has already been explored in the literature, for solving SAT and MaxSAT problems as QA systems scale (bian2020solving, ). The existing techniques allow encoding SAT and MaxSAT, with certain limitations, in an Ising problem compatible with sparse QA architecture. Is it possible to use more connected topologies? One may have larger QA hardware graphs with higher per-qubit connectivity and less separation between clusters (tiles) of qubits. New encoding strategies might be required as well because these changes will result in the ability to solve more difficult Ising problems. The encoding process will significantly improve by taking advantage of additional connectivity using new methods for problem decomposition, placing small Boolean functions, penalty modelling, etc.
Problem 33.
@@ 6.9. Quantum Merging Trees
6.10. Quantum Congested Clique Model
Problem 46.
@@ 6.11. Verification and Certification
Problem 50.
Consider the class $\text{QMA}^{+}(2)$ as the unentangled quantum proofs with non-negative amplitudes. The protocols in (10.1145/3564246.3585248, ) are examples of proof verification and property testing. Further, each protocol has a single isolated property testing task, relying on non-negative amplitudes. Can the protocols be generalized, so that one may transfer the results of (10.1145/3564246.3585248, ) to $\text{QMA}(2)$?
Problem 51.
@@
6.12. Quantum Machine Learning
Problem 56.
A hybrid algorithm already exists to solve medium and large instances of the multidimensional knapsack problem (garcia2021knn, ). Is it possible to explore using dynamic parameters instead of fixed parameters defined in (garcia2021knn, ) for the optimization process? The dynamic parameters might be capable of adjusting as the algorithm is executed, resulting in more powerful algorithms. Do reinforcement learning techniques help to adjust the parameters?
Problem 57.
@@ 6.13. Complexity
Problem 62.
We already know that
$\text{BPP}^{\text{QNC}_{\mathcal{O}(1)}}\subsetneq\text{QNC}^{\text{BPP}}$
$\text{QNC}^{\text{BPP}}_{\mathcal{O}(1)}\subsetneq\text{BPP}^{\text{QNC}}$
and
$\text{BPP}^{\text{QNC}^{\text{BPP}_{\mathcal{O}(1)}}}\subsetneq\text{BPP}^{\text{QNC}}\cup\text{QNC}^{BPP}$
are all separations of hybrid quantum depth, and are all according to search problems. Can the same separations be proved w.r.t. decision problems (arora2023quantum, )?
Problem 63.
Problem 64.
@@
Problem 65.
Problem 66.
Problem 67.
Problem 68.
Problem 69.
Problem 70.
Problem 71.
@@
Problem 72.
$\text{BPP}^{\text{QNC}_{\mathcal{O}(1)}}\subsetneq\text{QNC}^{\text{BPP}}$
$\text{QNC}^{\text{BPP}}_{\mathcal{O}(1)}\subsetneq\text{BPP}^{\text{QNC}}$
@@
$\text{BPP}^{\text{QNC}^{\text{BPP}_{\mathcal{O}(1)}}}\subsetneq\text{BPP}^{\text{QNC}}\cup\text{QNC}^{BPP}$
@@
Problem 75.
A quantum oracle has been developed to generate pseudorandom states, as reported in the literature (https://doi.org/10.4230/lipics.tqc.2021.2, ). A pertinent problem in this context is the unitary synthesis problem: whether it is possible to implement every $n$-qubit unitary transformation in $\text{poly}(n)$ time, relative to some classical oracle, remains a topic of active research and discussion.
Problem 76.
@@ 6.14. Classification
that is the set of all generalized Mycielski transformations of $K_{2}$ obtained from $k-2$ consecutive applications of $\mu_{r-1}(\cdot)$ possibly with different $r$s in every iteration. Let $\chi_{\text{loc}(\mathcal{G})}$ be:
$\chi_{\text{loc}(\mathcal{G})}=\min\{c\in\mathbb{N}\ |\ \exists\text{c-coloring with dim}(\mathcal{N})=1\}$
s.t. $\mathcal{N}$ is a finite von Neumann algebra. For every $G\in\mathcal{M}_{k}$, it holds that $\chi_{\text{loc}(\mathcal{G})}\geq k$. Let $\mathcal{K}_{n}$ be the quantum complete graph on $\text{Mat}_{n}$ equipped with the tracial $\delta\text{-form}\ \psi_{n}$. Define:
@@
Problem 82.
Is there a way to find the smallest graph (and the smallest number of colours) exhibiting a separation between classical and quantum chromatic numbers?
Problem 83.